At a price of $17.50 to play 18 holes on a golf course, 145 consumers pay to play a game of golf each day. Increasing the price to $22.00 causes the number of consumers to decline to 105.

What is the price elasticity of demand, expressed as a positive number? (Round you answer to two decimal places)
Answer #1

Price elasticity = % change in the demand / %change in the price.

% change in demand = 105-145 / 105+145

= -40 / 250

= 0.16

%change in price = 22 - 17.50 / 22 +17.50

= 4.5 / 39.50

0.11

Price elasticity = 1.45.